OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (1, −3) and has a slope of 2?

OpenStudy (imer):

Since its linear equation, it would be \[y−y0=m(x−x0)→y−(−3)=2(x−1)→y=2x+1\]

OpenStudy (anonymous):

But none of my answers look like that :( @imer

OpenStudy (anonymous):

2x − y = 5 2x − y = −5 2x + y = 5 2x + y = −5

OpenStudy (imer):

sorry, I forgot to carry the sign, \[y−y0=m(x−x0)→y−(−3)=2(x−1)→y=2x-5\]

OpenStudy (anonymous):

Wait so wich one is it..... 2x-y=5?

OpenStudy (imer):

First one; 2x − y = 5

OpenStudy (anonymous):

Thank you!!! Can you help with another problem?

OpenStudy (imer):

Sure just post it.

OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = the quantity of 3x minus 1, divided by 2, solve for g(f(3))

OpenStudy (imer):

\[f(x)=2x+1, g(x)=(\frac{ 3x-1 }{ 2 })\]\[gf(3)=\frac{ 3(2(3)+1)-1 }{ 2 } \rightarrow \frac{ 3(6+1) -1 }{ 2 } \rightarrow \frac{ 20 }{ 2 } \rightarrow 10\]

OpenStudy (anonymous):

Thank you very much.